OpenStudy (pokemon23):

solve for y: y/2 +c=d solve for y: y/2 +c=d @Mathematics

OpenStudy (pokemon23):

NJ

OpenStudy (anonymous):

subtract c, multiply by 2

OpenStudy (saifoo.khan):

Show all your working.

OpenStudy (anonymous):

show us your larks

OpenStudy (pokemon23):

yes saifoo :)

OpenStudy (saifoo.khan):

y/2 +c=d \[\frac y2 =d -c\] multiply 2 on both sides, \[y = 2(d-c)\] \[y = 2d-2c\]